A tie bar on a vertical pressing machine is 200cm long and 75mm square in cross section. Determine the stress and extension produced in the bar when it is subjected to an axial load of 250kN. Take E=2.1*10^5N/mm²


1
Expert's answer
2021-02-15T05:40:30-0500

E = stress / strain

Let F, L, A, ∆l be the force, length, area of cross section, and extension or respectively,

E = FL / A∆l

Stress = F x L = 250 000 N x 0.2 m = 50 000 N/m = 50 kN/m

∆l = FL / AE = (50 000 N/m) / (75 x 10-3 m)2(2.1x1011 N/m2) = 4.2x10-5 m = 0.042 mm